Constantinos M. Michael

3700 North Charles Street, Apt 710, Baltimore MD 21218

(443) 858-9725 • con@jhu.edu • http://ugrad.cs.jhu.edu/∼cmichae

Education The Johns Hopkins Univesity Baltimore, MD, U.S.A. • Bachelor of Science in Computer Sci., 2006 Cumulative GPA: 3.84/4.00 • Selected Coursework:

Computer Science Unix Programming · Data Structures · Automata & Computation · Computer System Fundamentals · Computer Networks · Info Retrieval and Web Agents · Operating Systems · Natural Lang Processing · Algorithms · Object Oriented Systems · Declarative Methods · Artificial Intelligence · Network Security · Databases

Mathematics Multivariable Calculus · Linear Algebra · Discrete Math · Math Models

Work History Bloomberg L.P. - Research and Development (2005) Software Engineer

• Designed and developed an SQL abstraction layer and command-line interface over the proprietary db.

• Established a library to enable integration and standardization of the Python language into the Bloomberg • Development Environment. Ported and rewrote legacy code in Python.

• Designed and developed a new standard for GUI specification and Client-Server interaction.

• Coded an automated data-miner capable of producing graphics-rich reports of product usage statistics.

The Johns Hopkins University - Department of Computer Science (2003) Head Course Assistant for 600.113 Internet

• Instructed hands-on web design skills, ranging from basic HTML, to Database-driven websites, to interactive content. Organized the course as a head Course Assistant, held office hours, assigned grades.

Cyprus National Guard - Information Technology Branch (during army service 2000-2002) Web Designer, Programmer, Network Administrator

• Designed and programmed the website for the Cyprus National Guard. Collaborated with a small team of both officers and drafted personnel for the website and other major projects.

• Headed the design and programming of a secure and efficient document retrieval server, client, and GUI.

• Drafted the security policies for numerous Army servers and administered the network infrastructure.

• Tutored web design (HTML and ASP) to officers. Took part in a team effort to promote Linux usage.

Prolix Corporation (incorporated 2001) 90% Owner, Programmer and Web Designer

• Created Prolix, a server side application for rapid deployment and management of database intensive websites, and used it to develop websites by contractual agreement.

• Retailed hardware and software after agreement with Multitech Corp., the largest regional importer.

• Collaborated with fellow Fulbright scholars to design a flexible media appliance.

University of Cyprus - MRDL (Multimedia Research and Development Laboratory) (1999) Web Designer

• Designed websites for the university research project intranet-extranet websites.

• Collaborated with the MRDL Team and proposed and presented changes to existing websites.

Hawker Website Creations (1999) 49% Partner, Programmer and Web Designer

• Designed a gaming community website for the leading Internet Service Provider in Cyprus, Cytanet.

• Created HawkerCart (flat DB and MySQL versions), applications for rapid creation of e-commerce sites.

MCM Developments (1995 - 1998) Computer-Aided Architectural Designer

• Designed a 150-house complex (Agros, Cyprus) and a multi-tower complex (Odessa, Ukraine.)

Skills

• Proficiency in software development in Java, Python, C, C++, C#, Perl, Assembly, Visual Basic.

• Familiarity with Regular Expressions, Threads, Sockets, Database Design, SQL, JSP, ASP, L^{A}TEX

• Familiarity with UNIX, GNU/Linux, Win32.

Honors and Awards

• Awarded the 2005 Computer Science Outstanding Undergraduate Award.

• Honored with the prestigious CASP Undergraduate Scholarship from the Cyprus Fulbright Commission.

• Awarded the First Prize in English Poetry from The English School, Nicosia, Cyprus.

• Included in the Dean’s List of Students, for all six completed semesters at the Johns Hopkins University.